OpenStudy (anonymous):

help please!!! In Angle ABC, C is a right angle. Find the remaining sides and angles. Round your answers to the nearest tenth. a = 3, c = 19 A. M

OpenStudy (anonymous):

@jim_thompson5910 can you help me with this?

jimthompson5910 (jim_thompson5910):

to find b, you use this equation a^2 + b^2 = c^2 so you would plug in a = 3 and c = 19, then solve for b

OpenStudy (anonymous):

is b 11.6?

jimthompson5910 (jim_thompson5910):

let me check one sec

jimthompson5910 (jim_thompson5910):

no it's a bit too small

OpenStudy (anonymous):

because i got 135 after doing 3^2=9 and 19^2 =144 then subtracted the two

OpenStudy (anonymous):

hmm i must have made a mistake somewhere then :/ let me check.

OpenStudy (anonymous):

i got 361 after doing 19^2 i see my mistake :)

jimthompson5910 (jim_thompson5910):

no worries, silly mistake

jimthompson5910 (jim_thompson5910):

so what do you get for b now?

OpenStudy (anonymous):

i got 18.7 this time.

OpenStudy (anonymous):

which would round to 18.8

jimthompson5910 (jim_thompson5910):

same here

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

angle A can be found by tan(A) = opp/adj tan(A) = 3/18.8 A = arctan(3/18.8)

OpenStudy (anonymous):

okay so the choices are between b and c

OpenStudy (anonymous):

because angle a cannot be larger than angle b correct?

jimthompson5910 (jim_thompson5910):

A is the smaller angle, yes

OpenStudy (anonymous):

im sorry i meant to say between A and C

OpenStudy (anonymous):

is the correct answer A?

jimthompson5910 (jim_thompson5910):

tan(B) = opp/adj tan(B) = 18.8/3 B = arctan(18.8/3) B = ??

OpenStudy (anonymous):

i dont understand any of that.

jimthompson5910 (jim_thompson5910):

do you have a calculator?

OpenStudy (anonymous):

yes , i do.

jimthompson5910 (jim_thompson5910):

and have you covered inverse trig functions?

jimthompson5910 (jim_thompson5910):

ok you would type in 2nd then the "tan" button then type in 3/18.8 and close off the parenthesis then hit enter

OpenStudy (anonymous):

no , this is supposed to be an algebra class but the work is trig and none of it has been covered.

jimthompson5910 (jim_thompson5910):

sounds like a very mixed up class to me

OpenStudy (anonymous):

i got 0.159?

OpenStudy (anonymous):

yeah i have complained several times but nothing was done about it.

jimthompson5910 (jim_thompson5910):

make sure you're in degree mode

jimthompson5910 (jim_thompson5910):

you are currently in radian mode at the moment

jimthompson5910 (jim_thompson5910):

i see, so they don't care that you haven't had the proper prerequisite classes at all? that's not good

OpenStudy (anonymous):

i don't know how to put my calculator in degree mode :/

OpenStudy (anonymous):

it says DEG at the top of my screen i'm guessing thats it?

jimthompson5910 (jim_thompson5910):

you have a TI calculator right?

jimthompson5910 (jim_thompson5910):

yes hit that button or option

OpenStudy (anonymous):

yes, TI-30Xa

jimthompson5910 (jim_thompson5910):

ok now try it again and tell me what you get

OpenStudy (anonymous):

i got the same thing ..

OpenStudy (anonymous):

it says it's in degree mode but keeps giving me that same answer for some reason

jimthompson5910 (jim_thompson5910):

hmm frustrating, idk why that's happening, i'm sry anyways, you can use google as a calculator if you type in "arctan(3/18.8) in degrees" without the quotes, you'll get this result http://www.google.com/search?hl=&q=arctan%283%2F18.8%29+in+degrees&sourceid=navclient-ff&rlz=1B3GGLL_enUS420US420&ie=UTF-8

jimthompson5910 (jim_thompson5910):

see how it says "arctan(3 / 18.8) = 9.06650274 degrees" so this rounds to 1 decimal place to get 9.1 degrees so A = 9.1 degrees

OpenStudy (anonymous):

so that would round to 9.1 :)

OpenStudy (anonymous):

so that means the answer is a then?

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

thank you!! and yes i understand that part of it now.

jimthompson5910 (jim_thompson5910):

ok that's great